Money Demand According to Liquidity Preference Theery, why is the Money Demand curve downwaed sloping? a because interest rates rise as the Bank of Canada reduces the quantity of money demanded b. because interest rates fall as the Bank of Canada reduces the Money Supply c because people will want to hold less money as the cost of doing so fals d. because people will want to hold more money as the cost of doing so falls Money Demand and Money Supply Model-Equilibrium Which of the following shifts Money Demand to the right? an increase in the Price Level b. a decrease in the Price Level c an increase in the interest rate d a decrease in income levels &. Aggregate Demand and Aggregate Supply Shifts If the stock market crashes, what would be the effect on Aggregate Dermand and how could the Bank of Canada offset those effects? a. Aggregate Demand increases by increasing the Money Supply b. Aggregate Demand increases; by decreasing the Money Supply c. Ageregate Demand decreases by increasing the Money Supply d. Aggregate Demand decreases; by decreasing the Money Supply 9. Marginal Propensity to Consume If the MPC- 5/8, what is the government purchases multiplier? a 1/6 b S/6 c. 6/s d. 6 0. Automatic Stabilizers re are automatic stabilizers but no deliberate action by policymakers, how would government expenditures and taxes change as output fals? a. Both povernment expenditures and taxes fal b. Both govemment expenditures and taxes rise c. Govermment expenditures rise and taxes fall d. Government expenditures fall and taxes rise
Answer #1

6. d is correct

Interest rate is the cost of holding money. It is the interest rate that can be earned on bonds. Thus when ut falls money demanded increases.

7. a is correct

Increase in price level means decrease in purchasing power of the dollar. To buy the same more money is needed which leads to increase in demand for money.

8. c is correct

Stock market crash reduces wealth which leads to decrease in consumption spending. This decreases Aggregate demand. Bank of Canada can offset this by increasing money supply.

9. d is correct

Multiplier= 1/(1-mpc)

= 1/(1-5/6)= 6